A hook is thrown from the edge of a fruit carrier with an initial horizontal velocity of 10 m/s. If the vessel is 20-m high, answer the following questions:

After how many seconds will the hook reach the water surface?
How far in the water will it reach?
1
Expert's answer
2018-11-14T14:56:09-0500

h=(gt^2)/2

t=√(2h/g)=√((2×20)/10)=2 s

The horizontal distance

L=vt=10×2=20 m